A couple of FEB’s 3rd year students decided to make some money by offering tutorial class to their juniors in first and second year. The seniors rented a room for RM300 for 3 hours from the faculty, and developed tutorial module for a few courses. The cost of printing the module handouts is RM5 each, and the tutor is paid RM25 per hour, for a total of RM75 for each tutorial session.



a. If students (junior) are charged RM20 per pax to attend the tutorial session, how many of them must enroll for the 3rd year students to break even?



b. If students (junior) are charged RM17 per pax to attend the tutorial session, how many of them must enroll for the 3rd year students to not having a loss?



c. A smaller room is available at the faculty for RM200 for 3 hours. The 3rd year students are considering this possibility. How would this affect the break even point? Use RM20 per pax as tutorial fee.


1
Expert's answer
2022-03-28T13:29:24-0400

Let x be the number of students (junior).


Total cost= 300+75+5x


a. Total revenue = 20x

Breakeven point : total cost=total revenue

300+75+5x=20x

375=15x

375/15=x

x=25


b. Total revenue = 17x

Breakeven point : total cost=total revenue

300+75+5x=17x

375=12x

375/12=x

x=31.25

x~32

32 students (junior) is needed so that the 3rd year students do not have a loss.


c. New total cost=200+75+5x

Total revenue=20x

Breakeven point : total cost=total revenue

200+75+5x=20x

275=15x

275/15=x

x=18.33

x ~ 19

The new cost is RM100 less than the previous cost, so the number of students (junior) needed to enrol becomes 19 students compared to 32 students. Less revenue is needed to cover the total cost.